OpenStudy (anonymous):

Solve: |3r - 12| = 0 Answer {-4, 4} {0, 4} {0, -4} {4}

OpenStudy (cwrw238):

which one of the options gives result of zero?

OpenStudy (cwrw238):

does r = -4 ?

OpenStudy (anonymous):

YEs?

OpenStudy (cwrw238):

mod(3(-4) - 12) = mod(-24) = 24

OpenStudy (cwrw238):

i used mod as i cant find the vertical lines on my keyboard

OpenStudy (anonymous):

its {4}

OpenStudy (anonymous):

do this: I 3r-12 I=0 +12 +12 I 3r I=12 /3 /3 ------>this will give you I r I =12/3 which is 4 so (r=4)